# Why Quantum-like Models of Cognition Work

### v1

@@ Abstract
It is shown that Brownian motions executed by state points of neural membranes generate a Schrödinger-like equation with $\hbar/m$ replaced by the coefficient of diffusion $\sigma$ of the substrates.
One of the fundamental problems of the entire body of pioneering work on quantum-like models of the social sciences (see Refs [1, 2, 3] and references therein) is that in spite of its successes, it is plagued by the absence of any empirical evidence of the quantum-like nature of the brain as well as by the interpretational problems associated with quantum mechanics. The brain is an open macroscopic hot and and noisy system in which decoherence should occur almost instantaneously. Nevertheless, there is strong evidence that classical probability theory fails to account for a number of observed psychological behaviour like the order effect [4, 5], the conjunction effect [6] and the disjunction effect [7, 8] which are all indicative of quantum-like behaviour.
The purpose of this paper is to show that it is precisely the stochastic nature of the neural processes underlying cognition that generates a Schrödinger-like equation with the factor $\hbar/m$ replaced by the square root $\sigma$ of the diffusion coefficient of the relevant neural substrates. This is in the spirit of Nelson [9, 10, 11] who showed that quantum mechanics is an emergent theory, resulting from underlying stochastic processes (Brownian motion) that are entirely classical. Similarly, the quantum-like behaviour of cognition can be traced to classical Brownian motion in neural substrates.
@@
There is a whole body of empirical and theoretical work which has established the stochastic behaviour of such substrates [12]. In particular, random walk and diffusion models for spike activity of a neuron have been proposed [13, 14]. Stated in the language of neurophysiology, the electrical state of polarization of the somatic and dendritic membrane may be represented by a “state point” which executes random walk. Each incoming elemental EPSP (Excitatory Post Synaptic Potential) moves the state point one unit toward the threshold of neuron firing, and each incoming elemental IPSP (Inhibitory Post Synaptic Potential) moves the state point one unit away from the threshold. If the average rate of incoming elemental EPSP and elemental IPSP are the same, there is an equal probability at any time that the state point moves either a unit toward or a unit away from the threshold, i.e. there is no “bias toward” either input. Immediately after the state point has attained the threshold and caused the production of an action potential, it returns to the resting potential, only to begin its random walk again. Such a model would be a simple random walk model.
@@
In reality, however, these two rates may sometimes be different, and in realistic physiological models it would be far more reasonable to assume that that there is some excess of either EPSP or IPSP input. In this case, the probability for the state point to move one unit toward the threshold will be different from the probability for it to move away from the threshold. Considered as a diffusion process, the difference between these probabilities can be considered a“drift velocity”, either toward or away from the threshold. It is this “random walk with drift” [13] which is of particular interest to quantum-like cognitive modelling.
Following Nelson, let us consider the following model.
(A) Let the “state point” of a neuron be specified by a single number $X(t)$, the membrane potential. As time passes and the electrical state of the membrane varies, the state point will move back and forth along a straight line, executing Brownian motion without friction. This means the ‘motion’ is a conservative diffusion process, i.e. a process in which there is on the average no exchange of energy between the state point and its neural background, i.e. the mutual exchanges of energy average out to zero but are responsible for “quantum-like fluctuations” of the system. The motion is described by a Markoff process in the state space.
To describe the spontaneous spiking of neurons, one must introduce a resting potential and an absorbing barrier, the threshold. If at any time the state point reaches the threshold, the neuron produces an action potential.
(B) The path of a state point driven by Brownian motion is written as the Itô stochastic differential equation (SDE)
$dX(t)=b_{f}(X(t),t)dt+\sigma dW_{f}(t)$ (1)
@@
where $b_{f}$ is the forward drift velocity (caused by EPSP) which depends on the current position $x=X(t)$, $\sigma$ is the square root of the diffusion coefficient and $dW_{f}(t)$ is a forward Wiener process. Since these processes are conservative, backward processes (caused by IPSP) exist,
@@
where $dW_{b}(t)$ is the backward Wiener process.
(C) The diffusion coefficient $\sigma^{2}$ is determined by the physiological characteristics of the membrane.
(D) The solutions $X(t)$ of the stochastic equations are known to be continuous at all state points but nowhere differentiable. Hence, Nelson suggested the following average forward and backward differentials which we adopt:
@@
where $E_{t}$ denotes the expectation conditional on $X(t)=x$. For differentiable curves $D_{f}X(t)=D_{b}X(t)=\dot{x}=v(t)$, the ‘velocity’ of the state point. Let us postulate Newton’s law for the stochastic acceleration,
@@
The forward and backward stochastic differential equations are equivalent to two Fokker-Planck equations
@@
where $\rho(x,t)$ is the probability density of the random variable $X(t)$. Addition of these two equations results in the continuity equation
@@
with the current velocity $v(x,t)=(b_{f}(x,t)+b_{b}(x,t))/2$. The difference of the forward and backward drifts $u(x,t)=(b_{f}(x,t)-b_{b}(x,t))/2$ is the osmotic velocity. Subtracting the two Fokker-Planck equations we get
@@
Using stochastic control theory, Guerra and Morato also derived the following differential equations for the functions $R$ and $S$:
@@
The first equation is the Hamilton-Jacobi-Bellman equation, i.e. the Hamilton-Jacobi equation with an additional stochastic term $V_{Q}$ which takes the form
@@
in terms of $\rho=e^{2R}$. It is the analog of the Bohm quantum potential. The second equation can also be written in terms of $\rho$ as
@@
by putting $\psi=\exp(R+iS)=\sqrt{\rho}\exp(iS/m\sigma)$ and separating the real and imaginary parts [15, 16]. The coefficient $\sigma$ clearly plays the role of the factor $\hbar/m$ in quantum mechanics.
@@
This is the ‘Nelson map’. It maps the probability distribution function and the current and osmotic velocities in the neural substrates to the wave function. In other words, it associates a diffusion process in the substrates to every solution of the Schröodinger-like equation (19).
It should be mentioned that entangled states exist in stochastic mechanics and that stochastic mechanics and quantum mechanics agree in predicting all the observed correlations at different times. The reader is referred to the papers by Faris [17] and Petroni and Morato [18] for details. It should also be pointed out that the violation of Bell’s locality is an intrinsically statistical phenomenon (it cannot produce superluminal signals!) and hence a stochastic framework should be the most appropriate for facing such a challenging problem.
Finally, it is well known that stochastic mechanics and Feynman’s path integral formulation of quantum mechanics are intimately related [19, 20]. Feynman’s path integral has been applied to finance [21] with some success. Hence a stochastic mechanics approach to quantum finance should also be possible.

### v2

# Rabindra Tirtha, New Town, Kolkata 700156, India

@@ Abstract
Despite the prevalent view that quantum mechanics is irrelevant to macroscopic biological systems because of inherent noise and decoherence, this paper demonstrates that the electrical noise (Brownian motion) in neuron membranes gives rise to an ‘emergent’ Schrödinger equation involving a new neuronal constant $\hat{\hbar}$, fundamentally challenging the standard view of neuronal behaviour. This result could provide new insights into the underlying mechanisms of brain function, thus challenging existing paradigms in both quantum physics and neuroscience. A possible empirical test of this emergent quantum behaviour would be to look for quantum fluctuations in subthreshold neural oscillations.
Quantum mechanics is widely regarded as the most fundamental theory of nature, suggesting its applicability to even brain functions. However, the prevailing view is that quantum coherence is lost at the neuronal level due to the large size and complex environment of the neurons, a process known as ‘decoherence’ [1, 2]. As a result, both physicists and neuroscientists have traditionally dismissed the relevance of quantum mechanics in neuronal processes. Unlike previous works that primarily focus on classical interpretations of stochastic processes in neurons, our approach explores the quantum-like dynamics that ‘emerges’ directly from such processes.
Building on Nelson’s method of deriving the Schrödinger equation from underlying classical stochastic processes [3] and its further developments [4, 5, 6, 7, 8, 9, 10], this paper demonstrates how stochastic neural dynamics can exhibit emergent quantum-like behaviour exhibiting coherence, stability and structure, a surprising example of ‘order-from-disorder’ [11]. Such a perspective offers a new framework for understanding brain function in terms of quantum-like processes.
Figure 1: Random walk with and without drift. The value $x(t)$ of the random variable $X(t)$ at time $t$ equals the last period’s value plus a constant (a drift) and a white noise. The drift can be in the direction of the threshold (forward) or away from it (backward).
Figure 2: Wiener process or gaussian random walk (Brownian motion) showing diffusion, i.e. increase of variance with time.
A substantial body of empirical and theoretical research has already established that neurons and neural networks exhibit stochastic behaviour [12, 13]. A stochastic process is a sequence of random variables whose values change over time in an uncertain way so that one knows only the distribution of the possible values at any point in time. It will be useful to introduce a few more related technical terms. A Markov stochastic process is a particular type of stochastic process where only the current value of a variable is relevant for predicting the future movement, not past values. A random walk is the stochastic process formed by successive summation of independent, identically distributed random variables. It is one of the most basic and well-studied topics in probability theory. A Wiener process, also called Brownian motion, is a Markov process which is essentially a series of normally distributed random variables such that for later times the variances of these normally distributed random variables increase, a process called ‘diffusion’.
Figure 3: Schematic representation of a neuron’s state evolving under stochastic dynamics. The figure illustrates the membrane potential’s Brownian motion, with the forward and backward processes corresponding to Excitatory (EPSP) and Inhibitory (IPSP) Post-Synaptic Potentials, respectively.
@@
Notably, random walk and diffusion models have been proposed to describe neuronal dynamics and spike activity. For instance, the electrical state of polarization in the somatic and dendritic membrane can be modeled as a ‘state point’ executing a random walk, influenced by excitatory and inhibitory post-synaptic potentials (EPSPs and IPSPs) [14, 15]. Each incoming elemental EPSP (Excitatory Post Synaptic Potential) moves the state point one unit toward the threshold of neuron firing, and each incoming elemental IPSP (Inhibitory Post Synaptic Potential) moves the state point one unit away from the threshold. If the average rate of incoming elemental EPSP and elemental IPSP are the same, there is an equal probability at any time that the state point moves either a unit toward or a unit away from the threshold, i.e. there is no “bias toward” either input. Immediately after the state point has attained the threshold and caused the production of an action potential, it returns to the resting potential, only to begin its random walk again. Such a model would be a simple random walk model.
@@
In reality, however, these two rates may sometimes be different, and in realistic physiological models it would be far more reasonable to assume that that there is some excess of either EPSP or IPSP input. In this case, the probability for the state point to move one unit toward the threshold will be different from the probability for it to move away from the threshold. Considered as a diffusion process, the difference between these probabilities can be considered a“drift velocity”, either toward or away from the threshold.
Our aim being to demonstrate that the inherent stochastic nature of neural processes, as specifically implied by ‘random walk with drift’ models, can generate a Schrödinger-like equation, we will treat the membrane potential $X(t)$ as a random variable and establish a connection between classical diffusion processes and emergent quantum-like behavior in neurons.
Following Nelson’s approach, let us then consider the following model. The “state point” $X(t)$ of a canonical neuron (consisting of a dendritic arbor, the soma and the axon) represents the membrane potential. As time passes and the electrical state of the membrane varies randomly, the state point moves back and forth along a straight line, executing Brownian motion without friction. Then $X(t)$ follows the stochastic differential equation (SDE)
$dX(t)=b_{f}(X(t),t)dt+\sigma dW_{f}(t).$ (1)
@@
Here, $b_{f}$ represents the forward drift velocity (caused by EPSPs), $\sigma$ is the square root of the diffusion coefficient, and $dWf(t)$ is a forward Wiener process. Since these processes are conservative, backward processes (caused by IPSP) also exist. The SDE for such backward processes is
@@
where $dW_{b}(t)$ is the backward Wiener process. The diffusion coefficient $\sigma^{2}$ is determined by the physiological characteristics of the membrane.
The solutions $X(t)$ of the stochastic equations are known to be continuous at all state points but nowhere differentiable. Hence, Nelson suggested the following average forward and backward differentials which we adopt:
@@
where $E_{t}$ denotes the expectation conditional on $X(t)=x$. For differentiable curves $D_{f}X(t)=D_{b}X(t)=\dot{x}=v(t)$, the ‘velocity’ of the state point.
To bridge this stochastic description with quantum mechanics, we now proceed to derive a Schrödinger-like equation from the underlying diffusion (Wiener) process. Following Nelson, let us postulate Newton’s law for the stochastic acceleration,
@@
As is well known, the forward and backward SDEs lead to two Fokker-Planck equations for the probability density $\rho(x,t)$ of the random variable executing Brownian motion:
@@
Adding these equations results in the continuity equation
@@
with the current velocity defined by $v(x,t)=(b_{f}(x,t)+b_{b}(x,t))/2$. This displays the role of the current velocity in maintaining the probability distribution:
The difference of the forward and backward drifts $u(x,t)=(b_{f}(x,t)-b_{b}(x,t))/2$ is defined as the osmotic velocity. Subtracting the two Fokker-Planck equations results in
@@
Using stochastic control theory and the current velocity as the control, Guerra and Morato showed that the following differential equations for the functions $R$ and $S$ extremize the action:
@@
The first equation is the Hamilton-Jacobi-Bellman equation, i.e., the Hamilton-Jacobi equation with an additional stochastic term $V_{Q}$ which takes the form
@@
in terms of $\rho=e^{2R}$. It is the analog of the Bohm quantum potential [16]. The second equation can also be written in terms of $\rho$ as
@@
by putting $\psi=\exp(R+iS/m\sigma)=\sqrt{\rho}\exp(iS/m\sigma)$ and separating the real and imaginary parts [7, 16]. The argument $x$ in the wave function $\psi(x,t)$ represents the ‘state point’ of the neuron membrane and plays the role of the spatial coordinate and the coefficient $\sigma$ plays the role of the factor $\hbar/m$ in standard quantum mechanics.
It should be borne in mind that the stochastic processes which occur in neural membranes are different from those that give rise to standard quantum mechanics which is responsible for the stability and structure of the atoms and molecules which constitute the neurons. It would therefore be convenient to introduce a new universal constant ^$\hat{\hbar}=m\sigma$ for neuronal media and rewrite the above equation in the form
$i\hat{\hbar}\frac{\partial}{\partial t}\psi(x,t)=\left(-\frac{\hat{\hbar}^{2}}{2m}\partial_{x}^{2}+V(x)\right)\psi(x,t).$ (20)
and treat it as the Schrödinger equation for canonical neurons. Like the standard Schrödinger equation in quantum mechanics, this equation will also ensure a novel level of stability and structure, i.e. stable order in the stochastic world of neurons.
@@
This is the ‘Nelson map’. It maps the probability distribution function and the current and osmotic velocities in the neurons to a wave function. In other words, it associates a diffusion process in the neurons to every solution of the Schröodinger-like equation (20).
Figure 4: Flow chart depicting the main mathematical steps used in deriving the Schrödinger equation from stochastic mechanics.
A possible experimental validation of the ideas proposed in this paper would be to look for ‘quantum-like fluctuations’ in subthreshold neural oscillations. Subthreshold oscillations are membrane oscillations that do not directly trigger an action potential because they do not reach the necessary threshold for firing. However, they may facilitate sensory signal processing [17]. Unlike oscillations that trigger action potentials and can be modelled by the charging and discharge of capacitors in RC circuits as in the Hodgkin-Huxley model [18, 19], subthreshold oscillations may be better modelled by LC circuits. The total energy of such a circuit is
$U=\frac{1}{2}(CV^{2}+LI^{2})=\frac{1}{2}\left(\frac{C^{2}V^{2}}{C}+\frac{L^{2}I^{2}}{L}\right)$ (24)
where $C$ is the capacitance, $L$ is the inductance, $V$ is the voltage and $I$ the current. This can be written in the form
$H$ $=$ $\frac{1}{2}\left(\frac{L^{2}I^{2}}{L}+\frac{C^{2}V^{2}}{C}\right)$ (25)
::=:= $\frac{p^{2}}{2m}+\frac{\omega^{2}Lx^{2}}{2}=\frac{p^{2}}{2m}+\frac{m\omega^{2}x^{2}}{2}$
with $U=H$ (the Hamiltonian), $L=m$ (the mass), $x=CV$ (the coordinate), $I=CdV/dt=dx/dt$ (the velocity), $p=LI$ (the momentum) and $\omega=1/\sqrt{LC}$ (the angular frequency). The identical formal structure of a mechanical and LC harmonic oscillator prompts us to speculate that the quantum mechanical description of the LC oscillator is in the form of a state vector $|\psi\rangle$. Then, in a representation in which $x$ is an independent variable and $|\psi(x,t)|^{2}dx$ is the probability that the oscillator has coordinate between $x$ and $x+dx$, the time evolution of the state function is governed by the equation of motion
$i\hat{\hbar}\partial_{t}|\psi\rangle=\hat{H}|\psi\rangle,\,\,\,\,\hat{H}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega^{2}\hat{x}^{2}.$ (26)
This corresponds exactly to the Schrödinger equation (20) with $\psi(x,t)=\langle x|\psi\rangle$, $\hat{p}=-i\hat{\hbar}\partial_{x}$, $V=m\omega^{2}x^{2}/2$. We therefore conclude that the Schrödinger equation (20) with $V=m\omega^{2}x^{2}/2$ describes the subthreshold oscillations of neurons.
Since we are interested in membrane oscillations, we need time-dependent solutions of (20) with $V=m\omega^{2}x^{2}/2$. Such states are called ‘coherent states’ $|\alpha\rangle$ of the harmonic oscillator. They were first introduced by Schrödinger in 1926, and their theory was developed further by Glauber [20] and Sudarshan [21] in connection with laser optics. Their importance lies in the fact that they are ‘minimum uncertainty’ states and closely approximate classical oscillators. One defines the ‘displacement operator’ $D(\alpha)$ as $D(\alpha)=e^{\alpha a^{\dagger}-\alpha^{*}a}$ where $\alpha=|\alpha|e^{i\phi}$ is a complex number. A coherent state $|\alpha\rangle$ is then defined as a displaced vacuum state $|0\rangle$,
$|\alpha\rangle$ $=$ $D(\alpha)|0\rangle,$ (27)
$a|0\rangle$ $=$ $0$ (28)
Using $\alpha(t)=\alpha e^{-i\omega t}=|\alpha|e^{-i(\omega t-\phi)}$, one can verify (see the Appendix for details) that the expectation value of $x$ in such a state is
$\langle x\rangle=\sqrt{2}x_{0}|\alpha|\cos(\omega t-\phi)$ (29)
where $x_{0}=\sqrt{\frac{\hat{\hbar}}{m\omega}}$ is the width of the harmonic oscillator ground state which is a gaussian distribution. The uncertainties in position and momentum in such a state satisfy the relation $(\sigma_{x})_{0}^{2}(\sigma_{p})_{0}^{2}=\hat{\hbar}^{2}/4$, showing they are minimum uncertainty states, i.e. states with minimum quantum uncertainty and hence closest to classical states.
Figure 5: Coherent state representation of the harmonic oscillator: the expectation value $\langle x\rangle$ as a function of time in the ground state of the harmonic oscillator, showing minimum uncertainty scatter points around a classical cosine wave.
Hence, the unequivocal prediction is that subthreshold neural oscillations should be sinusoidal with a scatter of ^$\hat{\hbar}^{2}/4$ around the classical values. A measurement of the scatter will therefore determine the value of $\hat{\hbar}$. Future work should therefore investigate quantum-like fluctuations in subthreshold neural oscillations using, for example, high-resolution electrophysiological recordings, such as patch-clamp techniques [22, 23], capable of detecting minute fluctuations in membrane potentials.
Another possibility is to search for the discrete stationary energy levels $E_{n}=(n+\frac{1}{2})\hat{\hbar}\omega$ of neurons implied by the harmonic oscillator model (26) (see the Appendix). These are solutions of the time-independent Schrödinger equation and should be observable during the quiescent or refractory periods of the neurons. The Helmholtz free energy $F$ and the average energy $\langle E\rangle$ of a quantum harmonic oscillator are given by (see the Appendix)
$F$ $=$ ^$\frac{\hat{\hbar}\omega}{2}+\frac{1}{\beta}\ln\left(1-e^{-\beta\hat{\hbar}\omega}\right),$ (30)
$\langle E\rangle$ $=$ ^$\frac{\hat{\hbar}\omega}{2}+\frac{\hat{\hbar}\omega e^{-\hat{\hbar}\omega}}{1-e^{-\beta\hat{\hbar}\omega}}$ (31)
where $\beta=1/kT$, $T$ being the absolute temperature. These expressions show that both $F$ and $\langle E\rangle$ tend to the zero-point energy as the temperature $T$ tends to zero. These results are experimentally verifiable in principle and can be used to determine the value of $\hat{\hbar}$.
Before concluding, it should be mentioned that ‘entangled states’ also exist in stochastic mechanics, and stochastic mechanics and quantum mechanics agree in predicting all observed correlations at different times. The reader is referred to the papers by Faris [8] and Petroni and Morato [9] for details. Looking for entanglement in neural systems is therefore another important area for further research, not only for its intrinsic value but also because entanglement is a key resource in quantum information processing, and it is important to find out whether the brain makes use of it, as conjectured [24]. We predict that the presence of non-classical correlations in neuronal fluctuations, akin to quantum entanglement, could be detected through cross-correlation analyses of simultaneous recordings from neighbouring neurons.
Finally, given that neural plasticity is often linked to the probabilistic nature of synaptic changes, our results suggest that quantum-like effects might influence how plasticity occurs in specific neural circuits. These insights could open new avenues for understanding cognitive processes, where classical and quantum descriptions might intersect, leading to a more comprehensive theory of brain function.
1 Appendix
When considering harmonic oscillators, it is convenient to introduce the ladder operators
$a$ $=$ $\frac{1}{\sqrt{2m\omega\hat{\hbar}}}(m\omega x+i\hat{p}),$ (32)
$a^{\dagger}$ $=$ $\frac{1}{\sqrt{2m\omega\hat{\hbar}}}(m\omega x-i\hat{p}).$ (33)
Using the commutation rule $[\hat{p},x]=-i\hat{\hbar}$, one gets the commutation relation $[a,a^{\dagger}]=1$, and
$a^{\dagger}|n\rangle$ $=$ $\sqrt{n+1}|n+1\rangle,$ (34)
$a|n\rangle$ $=$ $\sqrt{n}|n-1\rangle,$ (35)
$a^{\dagger}a|n\rangle$ $=$ $n|n\rangle$ (36)
where $n=0,1,2,...$ is an integer. The Hamiltonian operator can be written in the form
$\hat{H}=\hat{\hbar}\omega\left(a^{\dagger}a+\frac{1}{2}\right)=\hat{\hbar}\omega\left(N+\frac{1}{2}\right)$ (37)
where $N=a^{\dagger}a$ is the number operator. It then follows from the time-independent Schrödinger equation $\hat{H}\psi_{n}=E_{n}\psi_{n}$ that the energy eigenvalues are
$E_{n}=\hat{\hbar}\omega\left(n+\frac{1}{2}\right).$ (38)
The factor ^$\hat{\hbar}\omega/2$ is the ‘zero-point energy’ of neurons. The corresponding energy eigenfunctions are given by
$\psi_{n}(x)=\frac{1}{\sqrt{n!}}(a^{\dagger})^{n}\psi_{0}(x)$ (39)
with the ground state
$\psi_{0}(x)=\left(\frac{m\omega}{\pi\hat{\hat{\hbar}}}\right)e^{-\frac{m\omega x^{2}}{2\hat{\hat{\hbar}}}}$ (40)
which is a Gaussian distribution with width $x_{0}=\sqrt{\frac{\hat{\hbar}}{m\omega}}$.
Now note that the coherent state can be written as
$a|\alpha\rangle$ $=$ $\alpha|\alpha\rangle,$ (41)
$\langle\alpha|a^{\dagger}$ $=$ $\langle\alpha|\alpha^{*},$ (42)
$\langle\alpha|a^{\dagger}a|\alpha\rangle$ $=$ $|\alpha|^{2}.$ (43)
Using these results, one can compute the uncertainties in $x$ and $p$:
$(\sigma_{x})^{2}$ $=$ $\langle x^{2}\rangle-\langle x\rangle^{2}=x_{0}^{2}\left(n+\frac{1}{2}\right),$ (44)
$(\sigma_{p})^{2}$ $=$ $\langle p^{2}\rangle-\langle p\rangle^{2}=\frac{\hat{\hbar}^{2}}{x_{0}^{2}}\left(n+\frac{1}{2}\right),$ (45)
and hence for the ground state ($n=0$) $(\sigma_{x})_{0}^{2}(\sigma_{p})_{0}^{2}=\hat{\hbar}^{2}/4$. For further details of coherent states the reader is referred to Ref [25].
The thermodynamic properties of the quantum harmonic oscillator can be calculated using the standard techniques of statistical mechanics. The partition function
$Z=Tre^{-\beta\hat{H}}=\sum_{n=0}^{\infty}e^{-\beta E_{n}}=\sum_{n=0}^{\infty}e^{-\beta\left(n+\frac{1}{2}\right)\hat{\hbar}\omega}=\frac{1}{2}csch\left(\frac{\beta\hat{\hbar}\omega}{2}\right)$ (46)
where $\hat{H}$ is the Hamiltonian operator. Hence
$F$ $=$ $-\frac{1}{\beta}\ln Z=\frac{\hat{\hbar}\omega}{2}+\frac{1}{\beta}\ln\left(1-e^{-\beta\hat{\hbar}\omega}\right),$ (47)
$\langle E\rangle$ $=$ $-\frac{\partial\ln Z}{\partial\beta}=\frac{\hat{\hbar}\omega}{2}+\frac{\hat{\hbar}\omega e^{-\hat{\hbar}\omega}}{1-e^{-\beta\hat{\hbar}\omega}}.$ (48)